Determine all points of intersection
parabola y = x^2 + 3x -4 and line y = 5x + 11

Answer by HyperBrain(694)

We are given the following equations:
(1) y = x^2 + 3x -4
(2) y = 5x + 11
..
To find the intersections, equate (1) to (2).
x^2 + 3x -4 = 5x+11
x^2 -2x -15 =0
(x+3)(x-5) = 0
x= -3 and 5.
..
Using (2),
y=5(-3)+11=-15+11=-4
y=5(5)+11=25+11=36.
..
Thus, the points of intersection are (-3,-4) and (15,36).
Power up,
HyperBrain!